OpenStudy (anonymous):

How many mL would it take to prepare 500mL of a 0.10M Nitric Acid (HNO3) solution from a stock solution that is 69.0% pure and has a specific gravity of 1.42g/mL?

OpenStudy (anonymous):

I don't even have the slighest clue where to start on this one.

OpenStudy (anonymous):

ok i dont know how to go about specific gravity... but Im guessing you use that to calculate the molarity then you can simply do M1V1=M2V2 where you would plug the molarities and one volume to get the other.. hope that helps

OpenStudy (anonymous):

I think you have to use the purity and specific gravity to find the molarity, i'll see what i can come up with

OpenStudy (anonymous):

great good luck! and tell me what you find plz :)

OpenStudy (anonymous):

Well solution specific gravity is 1.42g/mL, 69.0% pure HNO3, so HNO3 specific gravity is .9798g/mL. Need molarity of HNO3 now

OpenStudy (anonymous):

Oh wait.. you can use the Molecular weight of HNO3 H=1 , N=14, O=16 (16x3=48) add them all up you get 63 g/mole and for molarity you need moles/litre so what you do here is .9798 g/ml / 63g/mole the g's will cancel out and you will get your molarity in moles/ml just divide that by 1000 to get it in litres and then you can solve the Q !

OpenStudy (anonymous):

so .9798 g/mL divided by 63g/mole = 64.29 moles/mL = .06429 Molarity? (moles/L) so then (500mL)/0.10M x (needed amt)/.06429M = 320.

OpenStudy (anonymous):

Sound close?

OpenStudy (anonymous):

ok.. when i divided .9798 by 63 i got 0.0156 moles/ml then divided that by a 1000 I got 1.56E-5 (0.0000156) then took that did M1V1=M2V2 M1V1/M2=V2 so (0.10M)(500)/1.56E-5 =3.205E6 ml but thats a very big number lol

OpenStudy (anonymous):

your answer however gives a reasonable amount... maybe I screwed up my calculations..0_o

OpenStudy (anonymous):

no depends on how you divided the .9798, i got 64.29

OpenStudy (anonymous):

i thought mine was incorrect, but your number is too big, mine is reasonable but i thought my math was wrong :)

OpenStudy (anonymous):

plus i think it's 320, then you add water to get 500mL